# EgoVLA: Learning Vision-Language-Action Models from Egocentric Human Videos

### v1

@@ Abstract
Real robot data collection for imitation learning has led to significant advances in robotic manipulation. However, the requirement for robot hardware in the process fundamentally constrains the scale of the data. In this paper, we explore training Vision-Language-Action (VLA) models using egocentric human videos. The benefit of using human videos is not only for their scale but more importantly for the richness of scenes and tasks. With a VLA trained on human video that predicts human wrist and hand actions, we can perform Inverse Kinematics and retargeting to convert the human actions to robot actions. We fine-tune the model using a few robot manipulation demonstrations to obtain the robot policy, namely EgoVLA. We propose a simulation benchmark called Isaac Humanoid Manipulation Benchmark, where we design diverse bimanual manipulation tasks with demonstrations. We fine-tune and evaluate EgoVLA with Isaac Humanoid Manipulation Benchmark and show significant improvements over baselines and ablate the importance of human data.Videos can be found on our website: https://rchalyang.github.io/EgoVLA/
@@ 1 Introduction
To benchmark robotic manipulation, we propose a new humanoid bimanual manipulation benchmark based on NVIDIA IsaacSim[17], namely Isaac Humanoid Manipulation Benchmark. In this benchmark, we provide 12 tasks including simple tasks that perform atomic actions and long-horizon tasks that combine multiple atomic actions. We collect 100 demonstrations per task and evaluate models using the benchmark. In our experiments, we first train our EgoVLA on our Ego-Centric Human Manipulation dataset and fine-tune on the humanoid manipulation demonstration collected for specific tasks. In our simulation benchmark, our EgoVLA outperforms different specialist and generalist baselines in both short-horizon and long-horizon tasks and achieves better generalization across visual observation and spatial locations.
@@ 3.3 Transferring EgoVLA to Humanoid Robot
4 Isaac Humanoid Manipulation Benchmark
@@
Beyond data scarcity, a major challenge in learning-based robotics is the lack of scalable, robust, and reproducible evaluation. Real-world evaluation is often costly, time-consuming, and raises concerns around safety and reproducibility—barriers that disproportionately affect resource-constrained settings such as academic labs. Recent work [69] has shown that simulation-based evaluations are highly correlated with real-world performance, supporting their use as a reliable proxy. To enable consistent benchmarking of humanoid manipulation, we introduce the Isaac Humanoid Manipulation Benchmark, built using NVIDIA Isaac Lab [70]. Isaac Humanoid Manipulation Benchmark is not designed for direct sim-to-real transfer. Instead, we leverage simulation, similar to LIBERO [71] and SIMPLER [69]—as a controlled and reproducible testbed for evaluating manipulation policies. Our simulated platform features the Unitree H1 [72] humanoid robot with two Inspire dexterous hands [73], and includes 12 manipulation tasks spanning both Short Horizon atomic actions (Push-Box, Flip-Mug, Pour-Balls, Close-Drawer, Open-Drawer, Open-Laptop, Stack-Can) and Long-Horizon, multi-stage skills (Sort-Cans, Insert-Cans, Unload-Cans, Insert-And-Unload-Cans, Stack-Can-Into-Drawer), as shown in Fig. 5.
@@ 5 Policy Trajectories Visualization
Figure 12: Robot Trajectories for Short-Horizon Tasks. We visualize the policy trajectories for short-horizon tasks in our Isaac Humanoid Manipulation Benchmark. The green dotted lines are the future robot hand trajectories predicted by our EgoVLA.
@@
Figure 13: Robot Trajectories for Long-Horizon Tasks. We visualize the policy trajectories for long-horizon tasks in our Isaac Humanoid Manipulation Benchmark. The green dotted lines are the future robot hand trajectories predicted by our EgoVLA.

### v2

@@ Abstract
Real robot data collection for imitation learning has led to significant advances in robotic manipulation. However, the requirement for robot hardware in the process fundamentally constrains the scale of the data. In this paper, we explore training Vision-Language-Action (VLA) models using egocentric human videos. The benefit of using human videos is not only for their scale but more importantly for the richness of scenes and tasks. With a VLA trained on human video that predicts human wrist and hand actions, we can perform Inverse Kinematics and retargeting to convert the human actions to robot actions. We fine-tune the model using a few robot manipulation demonstrations to obtain the robot policy, namely EgoVLA. We propose a simulation benchmark called Ego Humanoid Manipulation Benchmark, where we design diverse bimanual manipulation tasks with demonstrations. We fine-tune and evaluate EgoVLA with Ego Humanoid Manipulation Benchmark and show significant improvements over baselines and ablate the importance of human data.Videos can be found on our website: https://rchalyang.github.io/EgoVLA/
@@ 1 Introduction
To benchmark robotic manipulation, we propose a new humanoid bimanual manipulation benchmark based on NVIDIA IsaacSim[17], namely Ego Humanoid Manipulation Benchmark. In this benchmark, we provide 12 tasks including simple tasks that perform atomic actions and long-horizon tasks that combine multiple atomic actions. We collect 100 demonstrations per task and evaluate models using the benchmark. In our experiments, we first train our EgoVLA on our Ego-Centric Human Manipulation dataset and fine-tune on the humanoid manipulation demonstration collected for specific tasks. In our simulation benchmark, our EgoVLA outperforms different specialist and generalist baselines in both short-horizon and long-horizon tasks and achieves better generalization across visual observation and spatial locations.
@@ 3.3 Transferring EgoVLA to Humanoid Robot
4 Ego Humanoid Manipulation Benchmark
@@
Beyond data scarcity, a major challenge in learning-based robotics is the lack of scalable, robust, and reproducible evaluation. Real-world evaluation is often costly, time-consuming, and raises concerns around safety and reproducibility—barriers that disproportionately affect resource-constrained settings such as academic labs. Recent work [69] has shown that simulation-based evaluations are highly correlated with real-world performance, supporting their use as a reliable proxy. To enable consistent benchmarking of humanoid manipulation, we introduce the Ego Humanoid Manipulation Benchmark, built using NVIDIA Isaac Lab [70]. Ego Humanoid Manipulation Benchmark is not designed for direct sim-to-real transfer. Instead, we leverage simulation, similar to LIBERO [71] and SIMPLER [69]—as a controlled and reproducible testbed for evaluating manipulation policies. Our simulated platform features the Unitree H1 [72] humanoid robot with two Inspire dexterous hands [73], and includes 12 manipulation tasks spanning both Short Horizon atomic actions (Push-Box, Flip-Mug, Pour-Balls, Close-Drawer, Open-Drawer, Open-Laptop, Stack-Can) and Long-Horizon, multi-stage skills (Sort-Cans, Insert-Cans, Unload-Cans, Insert-And-Unload-Cans, Stack-Can-Into-Drawer), as shown in Fig. 5.
@@ 5 Policy Trajectories Visualization
Figure 12: Robot Trajectories for Short-Horizon Tasks. We visualize the policy trajectories for short-horizon tasks in our Ego Humanoid Manipulation Benchmark. The green dotted lines are the future robot hand trajectories predicted by our EgoVLA.
@@
Figure 13: Robot Trajectories for Long-Horizon Tasks. We visualize the policy trajectories for long-horizon tasks in our Ego Humanoid Manipulation Benchmark. The green dotted lines are the future robot hand trajectories predicted by our EgoVLA.